OpenStudy (anonymous):

Square tiles are used to cover the floor of a rectangular room. The rectangular room is 10 tiles long and 5 tiles wide. Find the number of tiles which touch the walls of the room.

OpenStudy (anonymous):

10*5=50 area of rectangle ,hcf ofarea=5 then area of square=5*5=25 then no of tiles=50/25=2

ganeshie8 (ganeshie8):

the question is about finding the number of tiles that touch the walls

ganeshie8 (ganeshie8):